# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ II-C The Description of Quantum Algorithm
Measurement (observation) is the only way to ascertain and extract the result of a quantum algorithm. From a physical implementation view, measurement embodies a degree of complexity. Observing or measuring implies an interaction between macroscopic or classical objects and a quantum system, with the process forming correlations and entanglements [37, 38]. Assume the state of the system under scrutiny to be: |$\ket{\phi}=\sum_{i}c_{i}\ket{i}$. The wave function of the entire system, together with the instrument, stands as following when the system is connected with the measuring apparatus:
|$\ket{\Psi}=\sum_{i}c_{i}\ket{i}\otimes e_{i}$ (32)
where {$e_{i}$} constitutes a set of the instrument’s wave functions. This correlation signifies a form of entanglement; once we detect that the instrument acquires the state |$\ket{e_{i}}$, the entire wave function collapses into the state |$\ket{i}\otimes\ket{e_{i}}$, and the system state |$\ket{i}$ can be inferred.
The physical realization of this process requires considerable experimental preparation, including tasks like calibrating the parameterized Hamiltonian, determining the ground state of the entire system, and so forth. Nonetheless, in this context, we focus solely on the algorithmic (circuit) level to utilize a simple example to illustrate the generation of entanglement, and the determining of the leftover segment when a partial state of an entangled system is observed.
Fig. 5 delineates the circuit employed to generate one of the Bell states - $\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. The first qubit |$\ket{0}$, transforms into a superposition state $\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ following its passage through the $H$ gate. Subsequently, through the CNOT gate, the |$\ket{1}$ within the first qubit triggers a flip in the associated |$\ket{0}$ of the second qubit, thereby creating an entangled state (eq. 33). Post the generation of entanglement, these qubits are unable to be expressed as a $\otimes$ product of independent subsystems. When an observation of the second qubit is made and it is located in |$\ket{0}$, the first qubit collapses into |$\ket{0}$. Conversely, if it is in the state |$\ket{1}$, the first qubit also becomes |$\ket{1}$. In this operation, the first qubit can be conceived as the quantum system requiring observation, whereas the second qubit stands as the measuring instrument.
Figure 5: The circuit to generate Bell state
@@
There are two possible outcomes of these observations, namely, $+1$ and $-1$, which correspond to the two eigenvalues of Pauli Z. The probability of obtaining $+1$ equals $|\alpha|^{2}$, whereas $|\beta|^{2}$ equates to the likelihood of garnering $-1$. The outcome of eq. 34 is the expected result of measurements. Specifically, if every observation yields a result of $+1$, it can be concluded that the qubit is in state |$\ket{0}$, and vice versa.
The aforementioned partial measurement on Bell state, whose observable can be written as $I\otimes Z$. Given a clear context, ’observed by $Z$’ can be simplified as measurements on the |$\ket{0}$ and |$\ket{1}$ basis (computational basis). It’s crucial to bear in mind the significance of the choice of the measurement basis. As an exemplification, for the state $\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$, using the |$\ket{0}$ and |$\ket{1}$ basis results in a $\frac{1}{2}$ probability for obtaining the post-measurement state |$\ket{0}$ and |$\ket{1}$, expressing the uncertainty of the result. However, if $\ket{+}$, $\ket{-}$ is employed (i.e., choosing observable Pauli $X$ for the measurement), the result $\ket{+}$ is guaranteed.
Being able to deterministically obtain the result of a quantum algorithm implies that the chosen measurement bases can reliably differentiate the final state. Although only orthogonal quantum states can assuredly be distinguished, the ultimate states of the quantum algorithm are occasionally non-orthogonal. In such instances, the objective is to ascertain the similarity rather than distinguishing the states reliably. This task can be proficiently executed by employing the subtle probability estimation of partial measurements. fig. 6 displays this technique, wherein the initial |$\ket{0}$ serves as an auxiliary qubit, transitioning into the superposition state $\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$ after passing through the left $H$ gate. Subsequently, under control of |$\ket{1}$ within this state, the circuit exchanges |$\ket{x}$ and |$\ket{y}$ to arrive at the final state eq. 35111The Swap gate can be implemented through the utilization of 3 CNOT gates.:
@@
Figure 6: swap test to evaluate the similarity of two states
If measuring auxiliary qubit alone by |$\ket{0}$ and |$\ket{1}$ bases, the likelihood of acquiring the post-measurement state |$\ket{0_{anx}}$ is:
$P(\ket{0_{anx}})=\frac{1}{2}+\frac{1}{2}|\bra{x}\ket{y}|^{2}$ (36)
$|\bra{x}\ket{y}|$ | is referred to as fidelity in the field of quantum information, fundamentally akin to the classical cosine distance. Obviously, the probability equals 1 if |$\ket{x}$ and |$\ket{y}$ are entirely overlapped, while the probability is $\frac{1}{2}$ if |$\ket{x}$ and |$\ket{y}$ are orthogonal. It’s vital to emphasize that probability estimation doesn’t hinge on the dimension. The larger the dimension of the state space, the greater the efficiency of this quantum trick. Inferred from this, the general similarity measure, $Euclidean\ distance=\sqrt{2-2|\bra{x}\ket{y}|}$, can also be computed. In practice, this trick forms the foundational basis for myriad quantum machine learning algorithms, inclusive of supervised and unsupervised machine learning [39], kernel approaches in quantum machine learning [40, 41], Quantum Support Vector Machine (QSVM) [42], Quantum Nearest-Neighbor Algorithms [43], and so forth.
@@ II-D The Description by Density Operators
The density operator is another more powerful mathematical language to describe quantum systems. The state vector can only represent pure states, while the density operator can represent pure and mixed states in a unified form. Besides, the density operator is a way more convenient tool to study quantum information, for example, the relationship between entanglement entropy and algorithm accuracy [44].
The density operator of a quantum system can be defined as follows. Suppose a quantum system is in pure state |$\ket{\psi_{i}}$ with probability $p_{i}$, where $i$ is the index of a set of state {|$\{\ket{\psi_{i}}\}$. Then, the quantum system can be described as an ensemble of pure states:
$\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ | (37)
A density operator is also called a density matrix, and generally, these two terms can be substituted for each other without distinction. If the density matrix $\rho$ satisfies (1) $tr(\rho)=1$; (2) $\rho$ is a positive semidefinite matrix, then $\rho$ is a legal representation of an ensemble $\{p_{i},\ket{\psi_{i}}\}$.
With the help of the density operator, we can distinguish pure and mixed states. The pure state refers to the quantum system being in a deterministic state $\rho=\ket{\psi}\bra{\psi}$ |; that is, the quantum system is in |$\ket{\psi}$ with probability 1. In all other cases, the quantum system is in a mixed state. The criterion to judge whether a quantum system is in a pure state or a mixed state is as follows: if $tr(\rho^{2})=1$, $\rho$ is a pure state; If $tr(\rho^{2})<1$, then it is a mixed state. According to density operators, the quantum algorithmic key steps are described as: 1) unitary evolution
$\rho_{2}=U\rho_{1}U^{\dagger}$ (38)
@@
$\begin{split}p(m)=&tr(M_{m}M_{m}^{\dagger}\rho)\\ \rho_{m}=&\frac{M_{m}\rho M_{m}^{\dagger}}{tr(M_{m}M_{m}^{\dagger}\rho)}\end{split}$ (39)
@@ II-E QC in Noisy Intermediate-scale Quantum Era
The quantum computer was conceptually put forward in the early 1980s [24]. QC hardware has made remarkable progress such that it is possible to deploy quantum algorithms in real. Noisy Intermediate-Scale Quantum (NISQ) computing devices [45] demonstrates the quantum superiority over classical computing [46]. We are not going to cover some well-known algorithms, such as Shor’s algorithm [31] and Grover’s algorithm [47], since they have been over-rendered by numerous materials [14]. Another primary reason we don’t introduce these algorithms is that the corresponding circuit depths to implement these algorithms are far beyond what is achievable in the NISQ devices. In the NISQ era, the number of operations (quantum gates) to implement a quantum algorithm should remain moderate to alleviate the suffering from decoherence and quantum gate errors.
In this section, we focus on the promising apps that are suitable for NISQ devices, including Variational Quantum Eigensolver (VQE) [48], QAOA [49], and some quantum machine learning algorithms. These algorithms have a common characteristic: using hybrid classical and quantum iterations to calibrate parameterized ansatzes. We will describe how these algorithms work and also discuss the challenge in this category at the end.
@@
VQE is mainly applied in computational chemistry. With the development of modern quantum chemistry, computational chemistry has become an important tool for drug synthesis, catalyst preparation, etc. However, chemical calculation involves a huge computation burden, whereas classical computers are stretched thin to provide adequate calculation accuracy and calculation scale. So, since quantum chemistry has to do with quantum systems, why not simulate computations on a quantum computer?
VQE simulates chemical computation on a NISQ quantum computer. It not only ensures the coherence of the quantum state but also achieves chemical accuracy for finding the ground state of a quantum chemical system. Its pipeline involves 3 major steps:
@@
$H=-\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}-\sum_{i}\frac{Ze^{2}}{r_{i}}+\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}}$ (40)
where $\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}$ is the kinetic energy term of the electron, $\sum_{i}\frac{Ze^{2}}{r_{i}}$ is the gravitational potential energy term between the nucleus and the electrons, and $\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}}$ is the repulsive energy term amongst electrons. This process is called the first quantization, and then $H$ in eq. 40 will be transformed to the structure represented by the fermion operators.
$H=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{r}a_{s}$ (41)
where $a_{j}^{\dagger}$ and $a_{j}$ are known as creation and annihilation operators, respectively. This process is called second quantization. After this process, the number of qubits to characterize $H$ will drop from scale $\mathcal{O}(mlog(n))$ to $\mathcal{O}(n)$ [50, 51], where $m$ and $n$ are the number of electrons and basis functions, respectively.
Finally, it needs to translate the fermionic Hamiltonian into Pauli operators. Fermionic Hamiltonian obeys the antisymmetry relationship, thereby the linear combination of translated Pauli strings should maintain this property. Exploiting Jordan-Wigner transformation [52], Parity transformation [53], Bravyi–Kitaev [53, 54, 55], or other improved methods [56], the $H$ composed by Pauli operators satisfying this criterion can eventually be obtained.
$H=\sum_{j}w_{j}H_{j}$ (42)
@@
The core step is sketched in Fig. 7. First, the trail state is initialized as the Hartree-Fock wavefunction. Simply speaking, for the molecular system with electronic spin-orbits, the Hartree-Fock state can be represented as |$\ket{0\dots 01\dots 1}$, where |$\ket{0}$ denotes empty orbit and |$\ket{1}$ denotes occupied orbit. Then, this state goes into the parameterized ansatzes (circuits) for evolving. All the parameters $\vec{\theta}$ (in Fig. 7) of the ansatzes could be initialized randomly or by some specific methods to alleviate the suffering from barren plateaus. When the trial wave function |$\ket{\varphi({\vec{\theta}})}$ is obtained, it must be rotated into the Z basis (computational basis) for measurements. In terms of the expectation value on each component (eq. 43), we can get the total expectation $\langle{H}\rangle$ by weighted summation eq. 42. Based on the $\langle{H}\rangle$ computed, we update the ansatz parameters using a certain optimization strategy and begin a new iteration until obtaining $\min\limits_{{{\vec{\theta}}}}\langle{H}\rangle$.
$\langle{H_{j}}\rangle=\bra{\varphi({\vec{\theta}})}U_{j}^{\dagger}(Z^{\otimes n})U_{j}\ket{\varphi({\vec{\theta}})}$ (43)
@@
The post-process, i.e., error mitigation, is an additional layer of processing on the measurement results aimed at reducing the impact of quantum noise for more accurate results. Since this thing is involved in all practical QC programs, we present a unified discussion after introducing other NISQ algorithms.
@@
The QAOA [49] is mainly used to solve combinatorial optimization problems. Due to the ubiquitousness of optimization in various scientific and technological domains, this algorithm has aroused broad research interests. The QAOA is an approximate version of the Quantum Adiabatic Algorithm (QAA) [57]. By applying the Suzuki-Trotter theorem, the time-dependent Hamiltonians of QAA are broken into a sequence of $2p$ time-independent Hamiltonians and induce two parameterized unitary operators $\hat{U}_{B}(\beta)=e^{-i\beta\hat{B}}$ and $\hat{U}_{C}(\gamma)=e^{-i\gamma\hat{C}}$, where $\hat{B}$ and $\hat{C}$ are the initial Hamiltonian and final Hamiltonian of QAA corresponding to “mixing Hamiltonian” and “problem Hamiltonian” of QAOA, respectively. Acting these two unitary operators with tuneable parameters $\beta$ and $\gamma$ on the initial state iteratively, the final state |$\ket{\beta,\gamma}$ will approximate the optimal solutions to the optimization problem.
|$\ket{\beta,\gamma}=\hat{U}_{B}(\beta_{p})\hat{U}_{C}(\gamma_{p})\dots\hat{U}_{B}(\beta_{1})\hat{U}_{C}(\gamma_{1})\ket{s},$ (44)
The schematic diagram of QAOA shown in fig. 8, it involves 3 major steps:
@@
In the vanilla QAOA for solving an unconstrained problem – MAXCUT, the initial state is prepared as $\ket{++\dots+}$, i.e., the uniform superposition over all possible states. When QAOA is applied to solve constrained problems, the straightforward way is to add penalty terms into the problem Hamiltonian to ensure infeasible solutions that violate the constraints are inferior to optimal feasible solutions when calculating the objective function. This scheme, say the Quadratic Unconstrained Binary Optimization (QUBO), does not modify the original operator framework, therefore can still use $\ket{++\dots+}$ as its initial state.
Despite this, numerous variant QAOA-based algorithms have shown that the choice of initial states can be more than $\ket{++\dots+}$. The naive approach is if a pre-processing method can obtain a biased initial solution (state) close to the optimal solution, then the convergence toward the final state should reach faster. This so-called “warm-starting” approach takes a wide variety of techniques, including variable relaxation, positive semi-definite programming, graph neural network, matrix product state, etc. to this end.
The other initial state preparation is for the alternative approach for handling constraints. That approach aims at restraining the evolution in the feasible subspace, thus $\ket{++\dots+}$ including infeasible states beyond the range of the feasible subspace is not applicable for the commencement. In this case, the initial state should be prepared as a superposition state composed of all or part of feasible solutions satisfying the constraints and excluding infeasible solutions. For a general constrained optimization problem, such as the Maximum Independent Set, preparing a uniform superposition state composed of all feasible solutions is equivalent to preparing an arbitrary uniform superposition state, which costs a lot. (exponential cost for deterministic manner [58] or post-measurement selection in a probability manner [59]). But this approach works particularly well for a certain class of NP-hard problems – Graph partition [60], maximum k-vertex cover [61], max-k colorable subgraph [62] etc., as their feasible subspace is spanned by all permutations of bit-strings with $Hamming\ weights=k$ (The number of $1s$ in a bit string is equal to $k$). That means their optima-searching space can be represented by a specific form of Dicke states – |$\ket{D^{n}_{k}}=\frac{1}{\sqrt{\tbinom{n}{k}}}\sum_{x\in\{0,1\}^{n},wt(x)=k}\ket{x}$, which can be easily prepared in deterministic manner [63].

### v3

@@ II-C The Description of Quantum Algorithm
Measurement (observation) is the only method to determine and extract the result of a quantum algorithm. Observing or measuring implies an interaction between macroscopic or classical objects and a quantum system, resulting in correlations and entanglements [34]. Assume the state of a system under observation is |$\ket{\phi}=\sum_{i}c_{i}\ket{i}$. When the system interacts with a measuring instrument, the combined wave function of the system and the instrument is given by eq. 34,
|$\ket{\Psi}=\sum_{i}c_{i}\ket{i}\otimes\ket{e_{i}},$ (34)
where |${\ket{e_{i}}}$ represents the set of wave functions of the measurement instrument. This correlation signifies a form of entanglement: once the instrument is observed to be in the state |$\ket{e_{i}}$, the entire wave function collapses to |$\ket{i}\otimes\ket{e_{i}}$, allowing the system state |$\ket{i}$ to be inferred.
The physical realization of this process involves significant experimental preparation, including tasks such as calibrating the parameterized Hamiltonian and determining the ground state of the entire system. However, in this context, we focus only on the algorithmic (circuit) level. We use a simple example to illustrate the generation of entanglement and the determination of the remaining segment when a partial state of an entangled system is observed.
Fig. 5 shows the circuit used to generate one of the Bell states: $\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. The process starts with the first qubit |$\ket{0}$, which is transformed into a superposition state $\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ after passing through gate $H$. Next, gate CNOT is applied, where |$\ket{1}$ in the first qubit flips |$\ket{0}$ in the second qubit, creating the entangled state in eq. 35. After generating entanglement, these qubits cannot be expressed as a $\otimes$ product of independent subsystems. When observing the second qubit, if it is found in state |$\ket{0}$, the first qubit collapses into |$\ket{0}$. Conversely, if the second qubit is in state |$\ket{1}$, the first qubit collapses into |$\ket{1}$. In this scenario, the first qubit can be viewed as the quantum system being observed, while the second qubit serves as the measuring instrument.
Figure 5: The circuit to generate Bell state.
@@
There are two possible outcomes when measuring the Pauli $Z$ observable: $+1$ and $-1$, corresponding to its eigenvalues. The probability of obtaining the result $+1$ is $|\alpha|^{2}$, while the probability of obtaining $-1$ is $|\beta|^{2}$. Thus, if every measurement yields $+1$, it can be inferred that the qubit is in the state |$\ket{0}$, and similarly, a result of $-1$ indicates the qubit is in the state |$\ket{1}$.
For a partial measurement of state Bell, where the observable is $I\otimes Z$, ‘observed by $Z$’ refers to measurements in bases |$\ket{0}$ and |$\ket{1}$ (the computational bases). The choice of measurement basis is crucial. For instance, for the state $\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$, measuring in the |$\ket{0}$ and |$\ket{1}$ basis yields a probability of $\frac{1}{2}$ for either outcome, reflecting the uncertainty of the result. However, if the measurement is performed using the $\ket{+}$ and $\ket{-}$ basis (i.e., measuring with the Pauli $X$ observable), the outcome $\ket{+}$ is certain.
Being able to deterministically obtain the result of a quantum algorithm implies that the chosen measurement bases can reliably differentiate the final state. Although only orthogonal quantum states can be distinguished with certainty, the ultimate states of a quantum algorithm are sometimes non-orthogonal. In such cases, the goal is to measure the similarity between states rather than distinguishing them with absolute certainty. This can be effectively achieved through probability estimation using partial measurements. Fig. 6 illustrates this technique, where an initial |$\ket{0}$ acts as an auxiliary qubit and transitions into superposition state $\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$ after passing through gate $H$. The circuit then uses state |$\ket{1}$ as control to perform a swap operation between |$\ket{x}$ and |$\ket{y}$, resulting in the final state given by eq. 37111The swap gate can be implemented using 3 CNOT gates.:
@@
Figure 6: Swap test to evaluate the similarity of two states.
If measuring the auxiliary qubit alone in bases |$\ket{0}$ and |$\ket{1}$, the probability of obtaining post-measurement state |$\ket{0_{anx}}$ is given by:
$P(\ket{0_{anx}})=\frac{1}{2}+\frac{1}{2}|\bra{x}\ket{y}|^{2},$ (38)
where $|\bra{x}\ket{y}|$ | is known as fidelity in quantum information theory, analogous to the classical cosine distance. This probability equals 1 if |$\ket{x}$ and |$\ket{y}$ are identical, and $\frac{1}{2}$ if |$\ket{x}$ and |$\ket{y}$ are orthogonal. It is crucial to note that this probability estimation is independent of the dimension of the state space; larger dimensions enhance the effectiveness of this quantum technique. Additionally, the general similarity measure can be computed as $Euclidean\ distance=\sqrt{2-2|\bra{x}\ket{y}|}$. In practice, this technique underpins many QML algorithms, including supervised and unsupervised ML, kernel methods in QML [35], quantum support vector machines (QSVM) [36], and quantum nearest-neighbor algorithms [37].
Lastly, it is worth mentioning weak measurement, which can be considered a generalization of strong (projective) measurement [38]. Unlike strong measurement, weak measurement does not directly observe the original system. Instead, it couples the system to an auxiliary quantum system, where partial information about the original system is extracted through a strong measurement of the auxiliary system, referred to as the probe. The wave function of the probe is typically a Gaussian wave packet centered at 0 with variance $\sigma^{2}$. Consider a simple example of a two-qubit coupling, where the wave function is described in eq. (39):
$\left(e^{-\frac{(x_{0}-\hbar/2)^{2}}{4\sigma^{2}}}\alpha\ket{0}+e^{-\frac{(x_{0}+\hbar/2)^{2}}{4\sigma^{2}}}\beta\ket{1}\right)\otimes\ket{x_{0}},$ (39)
where |$\ket{x_{0}}$ represents the post-measurement state of the probe system. When $x_{0}\approx 1$, the probability amplitude of the subsequent state vector |$\ket{0}$ is significantly high, and the converse is true when $x_{0}$ deviates from 1. If $\sigma\gg 1$, the measurement is considered very weak because the Gaussian term $e^{-\frac{(x_{0}-\hbar/2)^{2}}{4\sigma^{2}}}\approx 1$, indicating that the measurement induces minimal disturbance to the original system. In practice, the width of the probe’s wave packet, $\sigma$, is tunable; increasing $\sigma$ reduces the impact of the measurement on the original system. In essence, the state vector of the system gradually shifts in the direction of the probe’s collapse with a direct correspondence.
Weak measurement allows for information extraction from quantum systems while preserving quantum properties. Leveraging this capability, [39] introduced a technique that utilizes weak measurements to label additional samples, thus improving classification accuracy in active learning frameworks.
@@ II-D The Description by Density Operators
The density operator provides a more powerful mathematical framework for describing quantum systems compared to state vectors, which is limited to representing pure states. In contrast, the density operator unifies the description of both pure and mixed states, making it a more powerful tool in quantum mechanics. This operator is especially useful in quantum information theory, such as in analyzing the relationship between entanglement entropy and algorithmic accuracy [40].
The density operator for a quantum system is defined as follows: suppose a system is in a pure state |$\ket{\psi_{i}}$ with a probability $p_{i}$, where $i$ indexes a set of states |${\ket{\psi_{i}}}$. The quantum system is then described as an ensemble of pure states, with its density operator given by:
$\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}.$ | . (40)
The density operator is also referred to as the density matrix, and the terms are generally used interchangeably. For a density matrix $\rho$ to represent a valid quantum ensemble ${p_{i},\ket{\psi_{i}}}$, it must satisfy the following conditions: (1) $tr(\rho)=1$, and (2) $\rho$ is a positive semi-definite matrix. These properties ensure that $\rho$ is a legitimate description of a quantum system.
Using the density operator, we can effectively distinguish between pure and mixed quantum states. A pure state refers to a deterministic quantum system where $\rho=\ket{\psi}\bra{\psi}$ |, meaning the system is in state |$\ket{\psi}$ with probability 1. In contrast, a mixed state describes a quantum system that is in a probabilistic combination of different states. The criterion to differentiate between pure and mixed states is based on the trace of $\rho^{2}$: if $tr(\rho^{2})=1$, then $\rho$ is a pure state; if $tr(\rho^{2})<1$, the system is in a mixed state.
In terms of density operators, the quantum algorithmic core steps can be described as:
1) unitary evolution
$\rho_{2}=U\rho_{1}U^{\dagger},$ (41)
@@
$\begin{split}p(m)=&tr(M_{m}M_{m}^{\dagger}\rho)\\ \rho_{m}=&\frac{M_{m}\rho M_{m}^{\dagger}}{tr(M_{m}M_{m}^{\dagger}\rho)}\end{split},$ (42)
@@ II-E QC in Noisy Intermediate-scale Quantum Era
In this section, we do not cover widely known algorithms like Shor’s algorithm [29] and Grover’s algorithm [41], as they have been extensively explored in various materials [14]. Additionally, the circuit depths required to implement these algorithms far exceed what is currently feasible with NISQ devices. Instead, we focus on promising applications suited for NISQ devices, such as VQE [42], QAOA [43], and QML algorithms, which have demonstrated potential quantum superiority over classical computing [44]. These algorithms share a common feature: a hybrid classical and quantum iterations to calibrate parameterized ansatzes.
@@
VQE is mainly applied in computational chemistry. With the development of modern quantum chemistry, computational chemistry has become an important tool for drug synthesis, catalyst preparation, etc, which involves a huge computation burden. VQE simulates chemical computation on a NISQ quantum computer. It not only ensures the coherence of the quantum state but also achieves chemical accuracy for finding the ground state of a quantum chemical system. Its pipeline involves 3 major steps:
@@
$H=-\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}-\sum_{i}\frac{Ze^{2}}{r_{i}}+\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}},$ (43)
where $\frac{h^{2}}{2m_{e}}\sum_{i}\nabla_{i}^{2}$ is the kinetic energy term of the electron, $\sum_{i}\frac{Ze^{2}}{r_{i}}$ is the gravitational potential energy term between nucleus and electrons, and $\sum_{i}\sum_{j}\frac{e^{2}}{r_{ij}}$ is repulsive energy term among electrons. This process is called the first quantization, and then $H$ in eq. 43 is transformed to the structure represented by the fermion operators,
$H=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{r}a_{s},$ (44)
where $a_{j}^{\dagger}$ and $a_{j}$ are known as creation and annihilation operators, respectively. This process is called second quantization. After this process, the number of qubits to characterize $H$ drops from scale $\mathcal{O}(mlog(n))$ to $\mathcal{O}(n)$ [45], where $m$ and $n$ are the number of electrons and basis functions, respectively.
Finally, it needs to translate fermionic Hamiltonian into Pauli operators. Fermionic Hamiltonian follows the antisymmetry relationship, thereby the linear combination of translated Pauli strings should maintain this property. Exploiting Jordan-Wigner transformation [46], Parity transformation [47], or other methods [48], $H$ composed by Pauli operators satisfying this criterion can be obtained in eq. (45),
$H=\sum_{j}w_{j}H_{j},$ (45)
@@
The core step is sketched in Fig. 7. First, the trail state is initialized as Hartree-Fock wave function. Simply speaking, for the molecular system with electronic spin-orbits, the Hartree-Fock state can be represented as |$\ket{0\dots 01\dots 1}$, where |$\ket{0}$ denotes empty orbit and |$\ket{1}$ denotes occupied orbit. Then, this state goes into the parameterized devices (circuits) for evolving. All the parameters $\vec{\theta}$ of the devices is initialized randomly or by some specific methods to alleviate the suffering from barren plateaus in Fig. 7. When trial wave function |$\ket{\varphi({\vec{\theta}})}$ is obtained, it is rotated into basis Z (computational basis) for measurements. In terms of the expectation value on each component in eq. 46, we can get the total expectation $\langle{H}\rangle$ by weighted summation eq. 45. Based on $\langle{H}\rangle$ obtained, we update the device parameters using a certain optimization strategy and begin a new iteration until obtaining $\min\limits_{{{\vec{\theta}}}}\langle{H}\rangle$,
$\langle{H_{j}}\rangle=\bra{\varphi({\vec{\theta}})}U_{j}^{\dagger}(Z^{\otimes n})U_{j}\ket{\varphi({\vec{\theta}})}.$ (46)
@@
The post-process, i.e., error mitigation, is an additional layer of processing measurement results aimed at reducing the impact of quantum noise for more accurate results. Since it is involved in all practical QC programs, we present a unified discussion after introducing other NISQ algorithms.
@@
QAOA [43] is mainly used to solve combinatorial optimization problems. Due to the ubiquitousness of optimization in various scientific and technological domains, this algorithm has aroused broad research interests. QAOA is approximate to quantum adiabatic algorithm (QAA) [49]. By applying Trotter approximation, time-dependent Hamiltonians of QAA are broken into a sequence of $2p$ time-independent Hamiltonians. Two parameterized unitary operators $\hat{U}_{B}(\beta)=e^{-i\beta\hat{B}}$ and $\hat{U}_{C}(\gamma)=e^{-i\gamma\hat{C}}$ is induced, where $\hat{B}$ and $\hat{C}$ are initial Hamiltonian and final Hamiltonian corresponding to “mixing Hamiltonian” and “problem Hamiltonian” of QAOA, respectively. Acting two unitary operators with parameters $\beta$ and $\gamma$ on the initial state iteratively, final state |$\ket{\beta,\gamma}$ approximates optimal solutions for optimization problems in eq. (47),
|$\ket{\beta,\gamma}=\hat{U}_{B}(\beta_{p})\hat{U}_{C}(\gamma_{p})\dots\hat{U}_{B}(\beta_{1})\hat{U}_{C}(\gamma_{1})\ket{s}.$ (47)
QAOA schematic diagram involves 3 steps in Fig. 8:
@@
In the vanilla QAOA applied to solve an unconstrained problem such as MAXCUT, the initial state is typically prepared as $\ket{++\dots+}$, which is the uniform superposition over all possible states. For constrained problems, the straightforward approach is to incorporate penalty terms into the problem Hamiltonian. This ensures that infeasible solutions violating the constraints are inferior to optimal feasible solutions in terms of the objective function. This method, often referred to as quadratic unconstrained binary optimization (QUBO), does not change the operator framework, allowing the initial state to remain as $\ket{++\dots+}$.
Nevertheless, numerous QAOA variants have demonstrated that initial state choices need not be restricted to $\ket{++\dots+}$. A simple approach is to employ a pre-processing technique that yields a biased initial solution (state) closer to the optimal solution. This so-called ”warm-starting” approach accelerates convergence towards the final state. Various methods can be employed for warm-starting, including variable relaxation, positive semi-definite programming, graph neural networks, matrix product states, and more.
Another method for initial state preparation arises in an alternative approach to handling constraints, which focuses on restricting evolution to the feasible subspace. In this case, $\ket{++\dots+}$, which includes infeasible states, is not suitable. Instead, the initial state has to be prepared as a superposition composed of only feasible solutions that satisfy the constraints, excluding any infeasible solutions. For general constrained optimization problems, such as the maximum independent set, preparing a uniform superposition of all feasible solutions is computationally expensive [50]. However, this approach is particularly effective for a class of NP-hard problems, including graph partition [51], maximum k-vertex cover [52], and max-k colorable subgraph [53]. These problems have feasible subspaces spanned by permutations of bit-strings with specific Hamming weights (the number of 1s in a bit-string equals $k$). As a result, their search space can be represented by a specific form of Dicke states, |$\ket{D^{n}_{k}}=\frac{1}{\sqrt{\tbinom{n}{k}}}\sum_{x\in{0,1}^{n},wt(x)=k}\ket{x}$, which can be efficiently prepared [54].